OpenStudy (anonymous):

Jenny works at a video store.she earns $6.25 an hour plus $6 for each membership she sells. what is a formula? lets say a= total amount earned h= hours worked in one week b=number of memberships sold in a week

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

What do you think the equation is?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Let's fill in the variables with actual numbers first. Let's say that she works for 4 hours and she sells 2 memberships. How would you figure out how much she earned?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

add what together?

OpenStudy (anonymous):

4 & 2?

OpenStudy (anonymous):

Close but not quite. How much does she make every hour?

OpenStudy (anonymous):

i dont know it doesnt say but ummm 2?

OpenStudy (anonymous):

No, the original problem says that she make 6.25 for every hour. 4 & 2 are numbers I made up. If she works 4 hours and she makes 6.25 for every hour, what do you do with 6.25 and 4 to figure out how much she makes?

OpenStudy (anonymous):

* or +??

OpenStudy (anonymous):

Ok, well, what if she worked 100 hours. Would you multiply or add?

OpenStudy (anonymous):

m?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Ok, so we're going to multiply 6.25 by the number of hours this girl works. But is she getting paid only by the hour? Or is she getting extra money in another way?

OpenStudy (anonymous):

extra

OpenStudy (anonymous):

what extra money is she getting?

OpenStudy (anonymous):

$6

OpenStudy (anonymous):

$6 total? What if she sold 100 memberships? How much extra money would she get?

OpenStudy (anonymous):

106?

OpenStudy (anonymous):

Hmm, well it say that she will get $6 for EACH membership sold.

OpenStudy (anonymous):

oooh idk :/

OpenStudy (anonymous):

Well, here's a hint. When you see the word "each" in word problems, it usually means to multiply.

OpenStudy (anonymous):

So if I sold lemonade at $2 for each cup, and I sold 5 cups, I would make 5*2 = $10

OpenStudy (anonymous):

Did you understand the lemonade example?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Great. Understanding the words is a big part of writing the equation. Ok, so now we know that Jenny is making money 2 ways. Hourly and for each membership.

OpenStudy (anonymous):

Here's the problem again: Jenny works at a video store.she earns $6.25 an hour plus $6 for each membership she sells. what is a formula? lets say a= total amount earned h= hours worked in one week b=number of memberships sold in a week

OpenStudy (anonymous):

Now let's say that she worked 4 hours and sold 2 memberships again. Let's start with the hours. If she works 4 hours and gets paid $6.25 for every hour, what are you going to do with 4 and 6.25?

OpenStudy (anonymous):

m?

OpenStudy (anonymous):

Right. So now we have 4(6.25) for part of our equation. What about the extra money? How do we calculate that part? She sold 2 memberships and she get $6 for each membership. What are we going to do with 2 and 6?

OpenStudy (anonymous):

do you add the xtra money in? do we add? and add 2 and 6 into the equation?

OpenStudy (anonymous):

Well, we are going to add at the end.

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

But if she gets paid $6 for EACH membership sold and sold 2, how much extra money does she get?

OpenStudy (anonymous):

$12?

OpenStudy (anonymous):

Yes! What did you do with 6 and 2? +-* or /?

OpenStudy (anonymous):

umm multiply add or divide?

OpenStudy (anonymous):

well, does 6 + 2 = 12? Or does 6 * 2 = 12?

OpenStudy (anonymous):

so we multiply

OpenStudy (anonymous):

Yep. So now we have two parts. 6.25(4) for the hourly wages and 6(2) for the extra money. We are going to find the total amount that she makes. What should I do with the two parts? +-*/?

OpenStudy (anonymous):

make up the equation?

OpenStudy (anonymous):

yes. What are we going to do with 6.25(4) and 6(2)? add subtract multiply or divide?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

i mean multiply?

OpenStudy (anonymous):

Yes, we are going to add 6.25(4) and 6(2).

OpenStudy (anonymous):

that's how we're going to find the total amount of money that she makes!

OpenStudy (anonymous):

What did the 4 stand for again? What did the 2 stand for again?

OpenStudy (anonymous):

4 hours she work 2 memberships sold

OpenStudy (anonymous):

So now let's write it out 6.25(hours she worked) + 6(memberships sold) = total $$ Let's replace it with variables. I want you to try it.

OpenStudy (anonymous):

i honestly don't know

OpenStudy (anonymous):

a= total amount earned h= hours worked in one week b=number of memberships sold in a week

OpenStudy (anonymous):

i still don't know im being serious

OpenStudy (anonymous):

6.25(hours she worked) + 6(memberships sold) = total $$ Where do you think the 'a' will go? Well, I'll give you a hint. 6.25(hours she worked) + 6(memberships sold) = a I replaced a at the end because it stands for the total amount of money she made.

OpenStudy (anonymous):

Where should we place h and b?

OpenStudy (anonymous):

after 6.25

OpenStudy (anonymous):

which letter should go after 6.25?

OpenStudy (anonymous):

h

OpenStudy (anonymous):

Yes! So now we have: 6.25h + 6(memberships sold) = a

OpenStudy (anonymous):

you type the final equation.

OpenStudy (anonymous):

6.25 +6^2=a?

OpenStudy (anonymous):

What variable represents the number of memberships?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

not the value (number). the variable (the letter)... a= total amount earned h= hours worked in one week b=number of memberships sold in a week

OpenStudy (anonymous):

blm, you're almost there!!!! This is difficult to learn at first, so don't get discouraged.

OpenStudy (anonymous):

Ok, what I had you do to figure out the equation is to think about what you would do if you knew all of the numbers. So if she worked 4 hours and sold 2 memberships the equation would look like this: 6.25(4) + 6(2) = 37 But we DON'T know how many hours she worked, or the number of memberships so we can't figure out the total so we replace the numbers we don't know with variables. 4 = number of hours worked = h 2 = number of memberships = b 37 = total amount earned = a 6.25h + 6b = a That's your equation. Practice makes perfect. Ask your teacher for a list of key words. Plus usually means you have to add. Each usually means you have to multiply.

OpenStudy (anonymous):

thanks! :)